OpenStudy (anonymous):

write an explicit and a recursive formula for each sequence. 2,4,6,8,10

OpenStudy (anonymous):

aHH THE GOOD OLD 2+2+2+2+2 MAN HTIS BRIGNS BACK MEMORIES

OpenStudy (anonymous):

Recursive formula is easier to write, but not as useful, usually. To write the recursive formula, ask yourself "What is the rule for finding the NEXT term of the sequence?"

OpenStudy (anonymous):

2+2+2+2+2+2+2.... so \[A _{n} = A _{n-1}+2\]

OpenStudy (anonymous):

To find the explicit formula, think about what value you started at, what the difference is each time, and how many times you've added that difference. Here's an example sequence: 5,8,11,14,17 This one is +3,+3,+3,+3 and so on. To write the recursive formula would simple be \[A_{n+1} = A_n + 3\] However, what if I want to find the 10th term, but I don't want to use the recursive formula 9 times to get there? Well, I start at 5. I add 3 each time. If I'm at the 10th term, how many times have I added 3? Well, 9 times. So A10 = 5+10*3 and in general, I get \[A_n = 5+3(n-1)\] because 5 is where I start, 3 is the common difference, and n-1 will be how many times 3 gets added.